OBSERVATION OF TWO La(C_{3}H_{2}) ISOMERS FORMED BY DEHYDROGENATION OF PROPYNE

\underline{DILRUKSHI HEWAGE}, MOURAD ROUDJANE, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

C-H bond activation of small hydrocarbons is of importance in chemistry and industrial applications. La(C_{3}H_{2}) was formed by the reaction of laser-ablated La atoms and propyne (C_{3}H_{4}) in supersonic molecular beams. Two isomers of La(C_{3}H_{2}) were detected for the first time by mass-analyzed threshold ionization (MATI) spectroscopy. From the MATI spectra, the two isomers exhibit origin bands at 42953(5) and 43609(5) cm^{−1} and vibrational intervals of 425 and 535 cm^{−1}, respectively. They were identified as La(CCCH_{2}) formed from 1,3-dehydrogantion and La(HCCCH) formed by 3,3-dehydrogenation and were confirmed by measurements with deuterium substituted propyne (C_{3}D_{4}) as the precursor. The 1,3-dehydrogenated complex shows a higher ionization energy and larger metal-ligand stretching frequencies than the 3,3-dehydrogenated species. Based on DFT/B3LYP calculations, the electronic transitions responsible for the observed MATI spectrum of La(HCCCH)isomer is ^{1}A ← ^{2}A,and that of La(CCCH_{2})isomer is ^{1}A′ ← ^{2}A′.